OpenStudy (anonymous):

Given a right triangle, one leg is 6 cm and the hypotenuse is 10 cm. Use the Pythagorean theorem to find the length of the other leg. Please show ll work in order please dont explain I dont have time i have to turn it in i will give you a medal thanks!!!!

OpenStudy (whpalmer4):

 Pythagorean theorem says that in a right triangle, the sum of the squares of the lengths of the two legs equals the square of the length of the hypotenuse: \[a^2+b^2=c^2\]if you've labeled your triangle like I did. We know \(a = 6\) and \(c = 10\) so we can write\[6^2+b^2=10^2\]\[36+b^2=100\]Now just solve for \(b\). Subtract 36 from both sides, and take the square root of both sides to find the value of \(b\).

OpenStudy (whpalmer4):

it's far more important that you learn the concepts than that you get this problem turned in on time.

OpenStudy (anonymous):

i need to turn it in or i will fail

OpenStudy (whpalmer4):

Well, you'll probably fail sooner or later if you don't learn it. In any case, I laid out the steps for you.